OpenStudy (anonymous):

What is the restriction on the product of x^2+3x+ 2 over x^2-2x-3 multiplied by x^2+4x+3 over x+2

OpenStudy (anonymous):

\[\frac{ x^2 + 3x +2 }{ x^2-2x-3} * \frac{ x^2+4x+3 }{ x+2 }\] like this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{ (x+2)(x+1) }{ x-3)(x+1)} * \frac{ (x+3)(x+1) }{ x+2 }\] after factoring, cross out like terms, reduce /simplify.

OpenStudy (anonymous):

so x+1?

OpenStudy (anonymous):

you can't cross out x+3 and x-3

OpenStudy (anonymous):

oh true so (x+1)(x+3)/x-3?

OpenStudy (anonymous):

Yes, now restrictions: \[x \neq 3\] because it will be undefined at x=3 anything divide by 0 DNE

OpenStudy (anonymous):

Domain = ARN but x is not equal to 3

OpenStudy (anonymous):

thank you so much!